Tech Staff Meeting Notes July 7, 2009

The top priority is always repair of down stations.
NM problem stations
ET problem stations

The task list updated February 6, 2009 needs to be updated again. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


Priorities

I40 Weigh Station Install
AR Network survey
Ft Payne station
CVVA trigger upgrades
SM sample rate fix


News & Highlights

There was another fantastic Fiscal Fish Fry on June 26.

David is on the mandatory break in service for temporary employees. He'll be back on the payroll July 14.

Starting with this set of notes, we're including software and computer items. These will be discussed in more detail at the next meeting. These are currently listed under CERI node tasks. It isn't complete but its a start.


NMSZ

Short Period & Broadbands

Holly's NM list:

Down
MCAM.EHE
HALT

Muted Amplitude
GOBM.EHN
MARM.EHE
MARM.EHN
MLDM.EHE
TNMT.EHN
LRAR.EHE
CTCT.EHE

Isis Oxidation problem
SJBM.EHN
JHAR.EHE

Drifting
WYBT.EHZ
PPLM.EHZ

Gain Channel problem
MORT.EHG
CHRM.EHG

Unusable
CHRM
NNAR

Several of the stations can get improved telemetry by transmitting directly to the node. The original configuration changed enough so that there is room in the band. Jim will investigate.

NODE Tasks

MKTA: Keep getting tflr messages. David found a loose coax connection and replaced it. It works fine now. Al will return env_alarm.d to its original config (120 seconds I think for a MinOn though see the note Mitch left in the .d).

LNXT: Still need to tie generator status alarms into lnxtworm

NMAD: Repair barbed wire fence damaged by falling ice

AZTN: Upgrade PC and replace XLAK clock

CERI:

DSL is installed at lnxt and bosewell. Steve needs to complete the config at lnxt. Boswell needs the hardware installed. Telecom said we're too far from the switch to get it at mkta and nmad--we had ISDN at those sites so Steve is checking directly with local provider and, failing that, will investigate other options (e.g. satellite).

Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected).

ANSS SM

CVTN optical isolation. Need to upgrade to year 3 hardware.

RDST needs upgrade to year 3 hardware with router.

NAIT: Couldn't get the NWS cable modem to cooperate at NAIT (firewall?). Dean was going to ask Comcast about it and maybe get a new modem. Steve will follow up.

Need to complete firmware upgrade on CVVA. The facility changed hands and John is negotiating with the new owners to get back on the internet.

Now that CMG5 triggers actually contain data, we need to tune the trigger level. Mitch sent explanatory info and guidelines to Al and Steve to tune.

SLU had trouble getting triggers from the 10sps datastream. They are having success creating 200sps triggers off the 100sps channel without excessive triggers (though I wonder what their levels are). The two newer CMG5TD's we got from ASL also don't trigger well off the 10sps datastream. Jim will investigate the feasibility of switching to the 100sps tap as the trigger source for the 200sps upgrade. Then Al will retune as above.

Most of the strong motion stations can run at 100 sps without any modifications. We will install a router on the few that have problems (RDST will be the test case). These (eleven of them) will also need year 3 upgrades. ARRA funds have been requested for this.

HDBR
Recorders 1, 6, and 7 are in the shop for repair. Steve thinks they just need to be reformatted. Steve and David are working on getting them back in service so they can be resinstalled during the next lane closure. Al reports that of the remaining four, one is not associating.

HDAR
Need to remove concrete forms and configure everything to the correct ip addresses, etc to get the data flowing to Boswell and on to CERI. Have a strong signal from HDBT. Need to determine what import machine and wsv all these data will be on. How do we trigger.

HDBT
Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data.


ETSZ

Holly's ET list:
down
TVNC

Gain Channel problem
CCRT.EHG - is flat
SMNC.EHG- switches with SMNC.EHZ
ETT.EHG - is drifting

GMG.EHE - not sure- image attached

STAR Node

RCGA charging circuitry at is old and primitive and in need of an upgrade.

GMG - Need battery box solar reg mounting plate and need to fuse the battery.

BRBC Node

GFM - Fixed the bad channel but still needs some work to repair storm damage. The NC state park service is purchasing GFM. We already have a couple permissions from them and John initiated the paperwork for the GFM permission.

TRYN: the forest service wants to remove the tower. There will be a new 180' tower and we have permission to be on it. We also have permission to be in the building (rack mount only). Can we convert this station to AC?

SWET: Can get to the pc but not the sensor. John is investigating.

CPCT: The aluminum enclosure with the GPS breakout board at CPCT is corroded, fragile, and needs replaced. Need to re-work lighting and pest chaser. John has a pest chaser and will use some landscape or pool hardware for the lighting. A new breakout and transient suppression enclosure is ready. Needs a locking tower box. This will also be upgraded to the standard bb.

FPAL Installation is tentatively set for the week of August 3-7. John will confirm by July 24. Chris and David will assist.
Need to check with Horton on huddle test results to see if CMG6 #6240 has an odd resonance. Horton found a peak in the noise spectra that seemed to drift from 38Hz to 12Hz over the course of an hour-at least for the hour that he plotted though he did say it seems to be more the rule than the exception. This can't be natural and the data look good otherwise so the install is still go.
Jim installed the FPAL CMG3 in the test vault and the data look good. PDF's are at http://www.ceri.memphis.edu/people/mwithers/PDF/

GPS

James has a very nice SOH website at http://www.ceri.memphis.edu/people/jpdavis/gama/results.html.

Comm to McCarty is out. It appears to be an issue with the network at the school and a tech will be on site tomorrow to investigate.

Have permission to install freewaves at camp robinson so long as they don't interfere with anything currently operating (they shouldn't). Should be able to shoot to nearby EMA office--Ausbrooks is helping with the permission there.

Have a replacement antenna for RLAP from UNAVCO. If we can't break the choke ring free, we can just swap the electronics.

Troy (HCES) took a severe lightning hit or massive power surge. James replaced a bunch of stuff and got it going again. Mattioli co-located hardware is still fried.

Need to contact new owners (the water district) of the building at Miss. Co. Airport to refresh the permission. Chris contacted the county water dept will work with us any way they can. They also have internet on site there.

The next Davis brain dump is 7/22 and will include everything you ever wanted to know about the NetRS and about wireless topology and settings in the NM gps network.

AR Network

Three temp installs are done. Steve will get them in a wsv and on heli_ew.

Greg determined that current draw should be about 675mA at 12.0 VDC assuming a 70% transmit ratio for the freewave. Two 55-60 watt panels should suffice and to allow for temps down to 10 degrees F, we should use four 105ah batteries to get 14 days backup power.

FDL


MISCELLANEOUS

What to do with the Steiner brain dump DVD's?

Mitch made a seisnet group on umdrive. Go to http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password. There are several things on there including John's database. We can put contact info there too when its available. Only people in the seisnet group have access to this area. If you download something for update, don't forget to upload it again so that we all have the same version of the files.

The Oklahoma scream2scream connection doesn't seem to be as stable as we'd like so Steve is working with OGS to replace it with an ew export/import. The feed to NEIC is going but it would be good to verify with McManus at NEIC.

Rotational sensor is feeding an earthworm but still has IRIG crosstalk problem to fix. Its fine on an Oscope hooked to the output of the mux and looks fine in LabView. Its bad when viewed with wave_viewer. Need to see what the sac files look like to make sure it isn't just a wave_viewer feature.

Supply is checking the condition of Parker's old Suburban, Chris doubts it will run after a year sitting over there. He told her we would try to crank it up if needed. Looks like we could send old blue sometime soon. Maybe it would get on govdeals this year.

Pink Palace is working with Gary and Michelle on a new perc-like display area. We'll pull the receiver etc, when they start pulling out the old display (or sometime earlier if an opportunity arises).


CERI

Chris called P Plant electrical shop. They filled the generators soon after the outage. He said when there is a power failure as long as the generator is running that powers fuel tank pumps on campus they will bring fuel to keep us online. Chris will get all contact information posted near transfer switch and added to Jim's list.

3890 Power Chris is going to make a diagram for 3890. Good luck son.

3876 Power Chris is going to make a diagram for 3876. Good luck son.